$SPAD/src/input tschirn.input

The Axiom Team December 3, 2016

Abstract

Contents

2

— * —

draw(curve(3*(t**2-3),t*(t**2-3)),t = -3..3, [title "Tschirnhausen's Cubic"])

———-

References^{[1] nothing}

4